# Towards Physics-Guided Foundation Models

### v1

@@ Introduction
PGFM enhances foundation models by systematically integrating domain-specific physcial knowledge, improving performance, robustness, and interpretability for diverse applications. Future research in PGFM offers promising advancements in AI across fields like healthcare (Farhadloo et al. 2024b), geospatial analysis, and engineering. We plan to develop and evaluate a PGFM model, focusing on domain knowledge integration using metrics like performance sample complexity.
Advancing PGFM in future work will involve developing hybrid frameworks that combine the adaptability of retrieval-augmented generation (Lewis et al. 2020) with real-time updates to optimize domain adaptation and reduce training resource demands. We will also explore PGFMs in the context of recent multi-modal foundation models (Ravirathinam et al. 2024), focusing on integrating physical knowledge (Sharma et al. 2024; Sharma and Shekhar 2022; Sharma et al. 2022) from diverse input sources and its impact on training procedures and downstream applications. We will also explore additional criteria for comparing PGFMs with other machine learning models, including training methods and challenges such as model complexity, computational requirements, location dependency, and calibration needs (Farhadloo et al. 2024a, b). In the long term, we aim to address challenges such as explainability (Arrieta et al. 2020; Farhadloo et al. 2022), robustness, bias, data quality, and adaptability.

### v2

# Blue Sky Ideas: Towards Physics-Guided Foundation Models

@@ Introduction
PGFM enhances foundation models by systematically integrating domain-specific physcial knowledge, improving performance, robustness, and interpretability for diverse applications. Future research offers promising advancements in AI across fields like healthcare (Farhadloo et al. 2024b), geospatial analysis, and engineering. We plan to develop and evaluate a PGFM model, focusing on domain knowledge integration using metrics like sample complexity.
Advancing PGFM in future work will involve developing hybrid frameworks that combine the adaptability of retrieval-augmented generation (Lewis et al. 2020) with real-time updates to optimize domain adaptation and reduce training resource demands. We will also explore PGFMs in the context of recent multi-modal foundation models (Ravirathinam et al. 2024), focusing on integrating physical knowledge (Sharma et al. 2024; Sharma and Shekhar 2022; Sharma et al. 2022) from diverse input sources and its impact on training procedures and downstream applications. We will also explore additional criteria for comparing PGFMs with other machine learning models, including training methods and challenges such as model complexity, computational requirements, location dependency (Yang et al. 2025; Ghosh et al. 2024d, c, a, b, 2023, 2022), and calibration needs (Farhadloo et al. 2024a, b). In the long term, we aim to address challenges such as explainability (Arrieta et al. 2020; Farhadloo et al. 2022), robustness, bias, data quality, and adaptability.